OpenStudy (anonymous):

solve each the following 12X^2+7X=10 I got x= 0.66 x= -1.25 2. 1/x-1/x+4=1/3

OpenStudy (anonymous):

12x2+7x-10=0 ->x=-5/4,x=2/3

OpenStudy (anonymous):

which is what u got

OpenStudy (anonymous):

and how about the second one how do i do it